OpenStudy (anonymous):

Derivative help (x^2+2)^ln(cos(3x+1))

OpenStudy (anonymous):

THE LOG IS IN POWER.......

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

\[ \huge \frac{d}{dx} (x^2 + 2)^{Ln(\cos(3x + 1))}\]

OpenStudy (anonymous):

f = g(x)^h(x) ln f = h(x)ln(g(x)) df/fdx = h'(x)ln(g) + h(x)g'(x)/g(x)

OpenStudy (anonymous):

df/dx = f.(h'(x)ln(g) + h(x)g'(x)/g(x))

OpenStudy (anonymous):

Thanks for the steps